### v1

# primordial black holes; microlensing; dark matter

Abstract
Stellar microlensing is a powerful tool for probing dark matter in the form of planetary and stellar mass compact objects (COs), in particular primordial black holes (PBHs). Under standard assumptions, current observations exclude COs in the mass range $10^{-11}\lesssim M/M_{\odot}\lesssim 10^{4}$aking up all of the dark matter. We provide an overview, aimed at theorists working on PBHs, of the history, theory, observational status, and future prospects of the field.
keywords:
primordial black holes; microlensing; dark matter
[…]
1 Introduction and history
Stellar microlensing is the temporary magnification of a star which occurs when a compact object crosses the line of sight. It was first studied by Einstein, who concluded that ’there is no great chance of observing this phenomenon’ Einstein (1936). However in 1986 Paczyński showed that by monitoring millions of stars in the Magellanic Clouds (MC) it would be possible to probe compact objects (COs) in the Milky Way (MW) halo Paczyński (1986).
Various collaborations started microlensing surveys in the early 1990s, and the first MC candidate events were published in 1993 by EROS Aubourg et al. (1993), and MACHO Alcock et al. (1993). By the late 1990s the MACHO collaboration had observed significantly more events than expected from stars in the MC or MW disk. This indicated that a significant fraction of the Milky Way halo could be in the form of roughly Solar mass COs Alcock et al. (1997, 2000). Baryon budget arguments excluded such a large population of astrophysical objects (e.g. white dwarfs) Fields et al. (2000). Therefore it was suggested, e.g. Refs. Yokoyama (1997); Jedamzik (1998), that these events could be due to Primordial Black Holes (PBHs), black holes that may form in the early Universe from large overdensities Zel’dovich and Novikov (1967); Hawking (1971). However the EROS Ansari et al. (1996); Lasserre et al. (2000); Tisserand et al. (2007) and OGLE Wyrzykowski et al. (2009, 2011) collaborations did not see a significant excess of events, and instead constrained planetary and stellar mass COs to make up less than of order 10% of the MW halo.
In 2016 the discovery of gravitational waves from mergers of multi-Solar mass PBHs by LIGO-Virgo Abbott et al. (2016) generated a surge of interest in PBH dark matter (DM) Bird et al. (2016); Clesse and García-Bellido (2017); Sasaki et al. (2016); Carr et al. (2016). However microlensing observations over two decades by the OGLE collaboration have failed to find the large, $\gtrsim 10^{2}$, number of events expected (under standard assumptions) if a significant fraction of the MW halo is in the form of stellar mass CO Mróz et al. (2024). These long duration observations have also increased the largest CO mass probed to $\sim 10^{4}M_{\odot}$, while high cadence observations by Niikura et al. have decreased the lowest mass to $\sim 10^{-11}M_{\odot}$ Niikura et al. (2019).
We overview how stellar microlensing surveys probe the fraction of COs, such as PBHs, in the MW halo by observing stars in the MC or M31. Our aim is to provide an introduction to this field that is accessible to theorists who are interested PBHs, and specifically the possibility that planetary and stellar mass PBHs make up a significant fraction of the DM. First, in Sec. 2, we review the theory of stellar microlensing, before focusing on observations in Sec. 3. We then look at future prospects in Sec. 4, before concluding with a brief summary in Sec. 5.
For an overview of the history of PBHs, see Ref. Carr and Green (2025). Refs. Green and Kavanagh (2021); Carr et al. (2026) review the formation of PBHs and the full range of observational probes of their abundance. For a review of the theory of microlensing see, e.g., Ref. Mao (2012), while Refs. Moniez (2010); Jetzer (2014) overview the observational status as of the 2010s.
2 Theory
In this Section we overview the theory of stellar microlensing. First, in Sec. 2.1, we cover the basics of photometric microlensing, including individual events (Sec. 2.1.1) and the optical depth and event rate (Sec. 2.1.2). We then look at the differential event rate and parameter constraints (Sec. 2.2) and non-standard events (Sec. 2.3). Finally we conclude, in Sec. 2.4, with a brief overview of astrometric microlensing.
2.1 Basics
2.1.1 Individual events
In this subsection we overview the key properties of individual ’standard’ photometric microlensing events where a point source is lensed by a point lens. For more details see, e.g., the original microlensing paper by Paczyński Paczyński (1986) or the review by Mao Mao (2012).
[…]
Figure 1: The geometry of a stellar microlensing event. The left and right panels show the view perpendicular to and along the line of sight respectively. The solid black and dashed grey lines shows the trajectory of the compact object and the Einstein radius, $R_{\rm E}$, respectively. The orange star, black circle and blue square denote the source, the lens and the observer respectively. In the left panel the dotted green lines show the distance from the observer to the lens and source, $D_{\rm L}=xD_{\rm S}$ and $D_{\rm S}$ respectively. In the right panel the green dotted lines show the impact parameter, $b$, and its minimum value $b_{\rm min}$.
Microlensing occurs when the splitting of images caused by gravitational lensing is too small to be resolved, and instead the source is amplified by a factor Refsdal (1964); Paczyński (1986):
$A=\frac{u^{2}+2}{u\sqrt{u^{2}+4}}\,$, (1)
where $u=b/R_{E}$, $b$ is the distance of the lens from the line of sight to the source and $R_{\rm E}$ is the Einstein radius
$R_{{\rm E}}=2\left[\frac{GMD_{\rm L}(D_{\rm S}-D_{\rm L})}{c^{2}D_{\rm S}}\right]^{1/2}\,$, (2)
where $G$ is the Gravitational constant, $M$ is the mass of the CO, $D_{\rm S}$ is the distance from the observer to the source , and $D_{\rm L}$ is the distance to the lens. For a Large Magellanic Cloud (LMC) microlensing survey, $D_{\rm S}=49.6\,{\rm kpc}$ Pietrzyński et al. (2019) and
$R_{\rm E}\sim 10^{-4}\,{\rm pc}\sqrt{x(1-x)}\left(\frac{M}{M_{\odot}}\right)^{1/2}\left(\frac{D_{\rm s}}{50\,{\rm kpc}}\right)\,$, (3)
where $x=D_{\rm L}/D_{\rm S}$ is the fractional distance of the lens along the line of sight ($0\leq x\leq 1$).
A microlensing event is usually defined to occur when $u\leq 1$, which corresponds to $A\geq A_{\rm T}=1.34$. The timescale of the event, $\hat{t}$, is defined as the time taken to cross the Einstein diameter or radius. The EROS and OGLE collaborations use the Einstein radius crossing time, while the MACHO collaboration and this review use the Einstein diameter crossing time, $\hat{t}=2R_{\rm E}/v_{\rm t}$, where $v_{\rm t}$ is the transverse speed of the lens relative to the line of sight. For lensing of stars in the LMC by CO in the MW halo typically
$\hat{t}\sim 1\,{\rm yr}\,\sqrt{x(1-x)}\left(\frac{M}{M_{\odot}}\right)^{1/2}\,.$ (4)
Due to the mass dependence of the timescale, observations over a long duration are need to probe large masses, while high cadence observations (i.e. at short intervals) are required to probe small masses.
Because the timescale, $\hat{t}$, depends on the lens mass, distance and velocity, it is not possible to measure the lens mass from the timescale of a single standard micolensing event (this is often referred to as the microlensing degeneracy). However, as we will see in Sec. 2.2, the lens mass can be probed statistically from the distribution of timescales, given a model for the lens density and velocity distributions. The Einstein crossing time, $\hat{t}$, is different from the time for which the amplification exceeds the threshold value $A_{\rm T}=1.34$, $t_{\rm e}$, and has observational and theoretical advantages over $t_{\rm e}$  [^1] These advantages are well known in the observational microlensing community, but are hard to find original sources for.. Observationally, determining when the amplification crosses a particular threshold (and hence $t_{\rm e}$) is challenging, whereas $\hat{t}$ is found by fitting to the entire light curve. Theoretically, $t_{\rm e}$ depends on the impact parameter (as well as the lens mass, distance and velocity), which makes the calculation of the differential event rate more complicated.
[…]
Figure 2: The magnification factor, AA, Eq. (5), as a function of $2t/\hat{t}$ for $u_{\rm min}=b_{\rm min}/R_{\rm E}=0.25,0.5$ and $1.0$ (from top to bottom). The horizontal dotted black line shows the threshold for a microlensing event, $A_{\rm T}=1.34$.
The impact parameter relative to the Einstein radius, $u=b/R_{\rm E}$, can be written as
$u=\left(\frac{2t^{2}}{\hat{t}^{2}}+u_{\rm min}^{2}\right)^{1/2}\,$, (5)
where $u_{\rm min}=b_{\rm min}/R_{\rm E}$, $b_{\rm min}$ is the minimum value of the distance of the lens from the line of sight to the source, and time $t$ is defined so that $t=0$ when $b=b_{\rm min}$. Figure 1 shows the geometry of a microlensing event, as viewed parallel to and along the line of sight, including $R_{\rm E}$, $D_{\rm S}$, $D_{\rm L}$, $b$ and $b_{\rm min}$. Figure 2 shows the magnification factor $A$, given by Eq. (5), as a function of $2t/\hat{t}$ for different values of $u_{\rm min}$. This light curve, which is symmetric and achromatic (i.e. the same for all wavelengths), is often referred to as the Paczyński light curve.
2.1.2 Optical depth and event rate
The optical depth, $\tau$, is the probability that a source star is being lensed at a given instant, i.e. that it falls within the Einstein radius of a lens:
$\tau=\int_{0}^{D_{\rm S}}n(D_{\rm L})\pi R_{\rm E}^{2}(D_{\rm L})\,{\rm d}D_{\rm L}=4\pi G\int_{0}^{D_{\rm S}}\rho(D_{\rm L})\frac{D_{\rm L}(D_{\rm S}-D_{\rm L})}{D_{\rm S}}\,{\rm d}D_{\rm L}\,$, (6)
where $n(D_{\rm L})$ and $\rho(D_{\rm L})=n(D_{\rm L})/M$ are the number density and (mass) density of lenses respectively. Note that the dependence on the lens mass, $M$, of the number density and the square of the Einstein radius cancel, and hence the optical depth depends on the total mass density of lenses along the line of sight, but not their individual masses. For a very simple model of the Milky Way (MW) halo with constant density, and a circular speed $v_{\rm c}=\sqrt{GM(<D_{\rm S})/D_{\rm S}}\approx 220\,{\rm km\,s}^{-1}$, $\tau=v_{\rm c}^{2}/2c^{2}\sim 10^{-6}$ Mao (2012). This small optical depth means that to detect microlensing events it is necessary to monitor a large number, i.e. millions, of stars. Furthermore, it is extremely unlikely that a single star will undergo two independent microlensing events.
A more useful quantity is the rate, $\Gamma$}, at which events occur, and we summarize the calculation as outlined in Ref. Mao (2012). The area, ${\rm d}\tilde{A}$, swept out by a lens in a time interval, ${\rm d}t$, is given by the product of the diameter of the Einstein ring, $2R_{\rm E}$, and the distance traveled, $v_{\rm t}{\rm d}t$: $^{\rm d}\tilde{A}=2R_{\rm E}v_{\rm t}{\rm d}t=4R_{\rm E}^{2}{\rm d}t/\hat{t}$. The probability, ${\rm d}\tau$, of an individual star undergoing a new microlensing event is given by
${\rm d}\tau=\int_{0}^{D_{\rm S}}n(D_{\rm L})\,{\rm d}A\,{\rm d}D_{\rm L}=\int_{0}^{D_{\rm S}}n(D_{\rm L})\left(\frac{4R_{\rm E}^{2}}{\hat{t}}\right)\,{\rm d}t\,{\rm d}D_{\rm L}\,.$ (7)
The total number of new microlensing events in time ${\rm d}t$ is $N_{\star}{\rm d}\tau$, where $N_{\star}$ is the number of stars being monitored, and hence the event rate is
$\Gamma=N_{\star}\frac{{\rm d}\tau}{{\rm d}t}=N_{\star}\int_{0}^{D_{\rm S}}n(D_{\rm d})\left(\frac{4}{\pi\hat{t}}\right)\pi R_{\rm E}^{2}{\rm d}D_{\rm d}=\frac{4N_{\star}}{\pi}\int\frac{{\rm d}\tau}{\hat{t}}\,$. (8)
Assuming that all events have the same timescale $\hat{t}$, leads to a rough relationship between the event rate and optical depth:
$\Gamma\sim\frac{4N_{\star}}{\pi\hat{t}}\tau\,.$ (9)
2.2 Differential event rate and parameter constraints
Since the timescale of a microlensing event depends on the CO mass (as well as its position and velocity), the differential event rate, $^{\rm d}\Gamma/{\rm d}\hat{t}$, i.e. the rate at which events of timescale $\hat{t}$, occur can be used to probe the CO mass. The standard expression for the LMC differential event rate, for a halo composed entirely of COs ($f=1$, where $f$ is the fraction of the total mass of the halo in CO) with mass $M$ and a smooth density distribution $\rho(r)$ is Alcock et al. (1997):
$\frac{{\rm d}\Gamma_{\rm MW}}{{\rm d}\hat{t}}=\frac{32D_{\rm S}}{M{\hat{t}}^{4}v_{{\rm c}}^{2}}\int^{1}_{0}\rho(x)R^{4}_{{\rm E}}(x)\exp{\left[-Q(x)\right]}{\rm d}x\,$, (10)
where $Q(x)=4R^{2}_{{\rm E}}(x)/(\hat{t}^{2}v_{{\rm c}}^{2})$. Note that Eq. (10) assumes that the velocity distribution, $f({\bf v})$, is Maxwellian
$f({\bf v})\,{\rm d}^{3}{\bf v}=\frac{1}{\pi^{3/2}v_{\rm c}^{3}}\exp{\left(-\frac{v^{2}}{v_{\rm c}^{2}}\right)}\,{\rm d}^{3}{\bf v}\,$, (11)
however this assumption is only exact if the density varies with Galactocentric radius $r$ as $\rho(r)\propto r^{-2}$ for all $r$.
[…]
Figure 3: The differential event rate, $^{\rm d}\Gamma/{\rm d}\hat{t}$, as a function of Einstein diameter crossing time, $\hat{t}$, for a standard halo (see text for details) composed entirely, $f=1$, of COs with mass $M=1M_{\odot}$ as given by Eq. (13).
The standard halo model historically assumed in microlensing studies (‘Model S’ in e.g. Ref. Alcock et al. (1997)) is a cored isothermal sphere with a density profile
$\rho(r)=\rho_{0}\,\frac{r_{{\rm c}}^{2}+r_{0}^{2}}{r_{{\rm c}}^{2}+r^{2}}\,$, (12)
and local dark matter density $\rho_{0}\equiv\rho(r_{0})=0.0079M_{\odot}{\rm pc}^{-3}$, core radius $r_{{\rm c}}=5$ kpc, Solar radius $r_{0}=8.5$ kpc and circular speed $v_{\rm c}=220\,{\rm km\,s}^{-1}$  [^2] There have been changes to some of best fit parameter values, for instance the Solar radius has been measured as $r_{0}=(8.18\pm 0.01\pm 0.02)$ kpc by the GRAVITY collaboration GRAVITY Collaboration et al. (2019). However, these changes have a smaller effect on the differential event rate than uncertainties in the density profile.. For the standard halo model, the differential event rate, Eq. (10), becomes Alcock et al. (1997)
$\frac{{\rm d}\Gamma_{\rm MW}}{{\rm d}\hat{t}}=\frac{512\rho_{0}(R_{{\rm c}}^{2}+R_{0}^{2})D_{\rm S}G^{2}M}{{\hat{t}}^{4}{v_{{\rm c}}}^{2}c^{4}}\int^{1}_{0}\frac{x^{2}(1-x)^{2}}{A+Bx+x^{2}}\exp{\left[-Q(x)\right]}{\rm d}x\,$, (13)
where $A=(r^{2}_{{\rm c}}+r^{2}_{0})/D_{\rm S}^{2}$, $B=-2(r_{0}/D_{\rm S})\cos{b}\cos{l}$ and $b=-32.8^{\circ}$ and $l=281^{\circ}$ are the Galactic latitude and longitude, respectively, of the LMC. Figure 3 shows the differential event rate for a standard halo composed entirely, $f=1$, of COs with $M=1M_{\odot}$. For different masses, the differential event rate scales as $M^{-1/2}$ and the timescales as $M^{1/2}$ Griest (1991). For short-timescale events ${\rm d}\Gamma/{\rm d}\hat{t}\propto\hat{t}^{2}$ while for long-timescale events ${\rm d}\Gamma/{\rm d}\hat{t}\propto\hat{t}^{-4}$, independent of the CO distribution Alcock et al. (1996); Zhao et al. (1996); Mao and Paczyński (1996).
Sahu pointed out that stars ($\star$’) in the LMC can act as lenses Sahu (1994), a phenomenon often referred to as ’self-lensing’. There are also contributions to the total differential event rate from stars in the MW disc and and COs in the LMC:
$\frac{{\rm d}\Gamma}{{\rm d}\hat{t}}=f\frac{{\rm d}\Gamma_{\rm MW}}{{\rm d}\hat{t}}+\frac{{\rm d}\Gamma_{{\rm MW},\star}}{{\rm d}\hat{t}}+\frac{{\rm d}\Gamma_{{\rm LMC},\star}}{{\rm d}\hat{t}}+f\frac{{\rm d}\Gamma_{\rm LMC}}{{\rm d}\hat{t}}\,$. (14)
The self-lensing rate is largest for observing fields in the centre of the LMC, where the number of stars is highest. The MACHO collaboration found, using models from Ref. Gyuk et al. (2000), an optical depth $\tau_{{\rm LMC},\star}\sim(1-3)\times 10^{-8}$. Using the same LMC model with some minor modifications, OGLE III and IV (which has a wider observing field) found 5.7 events expected from stars in the LMC (compared with the $10^{2-3}$ expected if all of the halo is in planetary or stellar mass COs) Mróz et al. (2024). The MACHO collaboration found an optical depth for microlensing by stars in the MW disc, $\tau_{{\rm MW},\star}\lesssim 10^{-8}$, lower than their estimate for self-lensing Alcock et al. (2000). However more recent studies have found significantly larger MW disc rates, comparable with, or even larger, than the self-lensing rate, in part due to a larger thick disc contribution Calchi Novati and Mancini (2011); Mróz et al. (2024). For the two MW models they considered Han and Gould (2003); Cautun et al. (2020), OGLE III and IV found 7.0 and 14.7 expected events respectively from stars in the MW disc Mróz et al. (2024). Note, however, that the relative numbers of events from different lens populations depend on the fields observed, and the detection efficiency (see below).
The total expected number of events, $N_{\rm exp}$, is given by
$N_{\rm exp}=E\int_{0}^{\infty}\frac{{\rm d}\Gamma}{{\rm d}\hat{t}}\epsilon(\hat{t})\,{\rm d}\hat{t}\,$, (15)
[…]
${\cal L}(f,M)=\exp{(-N_{\rm exp})}E\prod_{i=1}^{N_{\rm obs}}\frac{{\rm d}\Gamma}{{\rm d}\hat{t}}(\hat{t}_{i})\epsilon(\hat{t}_{i})\,$, (16)
where $N_{\rm obs}$ is the number of observed events and $\hat{t}_{i}$ are their timescales.
The MACHO first year results paper Alcock et al. (1997) studied a very wide range of models for the MW halo, some of which are not consistent with current observations. They found that there was a significant variation in the differential event rate, and hence the observational constraints on the CO mass and halo fraction. The OGLE collaboration finds that the uncertainty in their constraint is relatively small (see Extended Data Fig. 1 of Ref. Mróz et al. (2024)), however they only consider two benchmark models (with fixed parameter values) and assume a Maxwellian velocity distribution. Other studies Hawkins (2015); Green (2017); Calcino et al. (2018); Garcia-Bellido and Hawkins (2024) have found that the uncertainty in exclusion limits can be significant. Ref. Green (2025) found that the rate of long-timescale events (and hence the constraints on low-mass WIMPs) has a fairly weak dependence on the CO density and velocity distribution. However the rate of short-timescale events (and hence the constraints on high-mass WIMPs) depends sensitively on the local DM velocity distribution. An accurate calculate of the differential event rate (and parameter constraints) for models consistent with current observations, and numerical simulations of Milky Way-like galaxies, is an outstanding issue.
Microlensing observations towards the Small Magellanic Cloud (SMC) could in principle help discriminate halo-lensing from self-lensing, as the visible components of the SMC are elongated along the line of sight. The self-lensing rate is therefore expected to be larger than for the LMC. Furthermore for halo lenses, the timescales of events towards the LMC and SMC should be similar whereas for self-lensing, events towards the SMC are expected to be longer (due to the larger lens-source separation, and hence Einstein radius) Sahu and Sahu (1998). There are, however, significant uncertainties in the SMC’s structure due to its tidal interactions with the LMC. See Refs. Palanque-Delabrouille (1998); Graff and Gardiner (1999) for discussion of the status as of the late 1990s in the context of microlensing, and, e.g., the Introduction of Ref. Povick et al. (2025) for the current status in a more general context.
Crotts explored the advantages and disadvantages of microlensing observations towards M31 (our nearest neighbour galaxy, also known as Andromeda) Crotts (1992). In particular he pointed out that the high inclination of M31’s disc with respect to the line of sight would lead to a large asymmetry in the microlensing rate from lenses in the halo of M31. The event rate for the far side of the disc would be much higher than that for the near side, due to the larger column density of halo lenses. This would be a potentially powerful tool for discriminating events due to halo lenses from self-lensing or variable stars. Baillon et al. Baillon et al. (1993) pointed out that the large number of source stars is another advantage of M31 over the MCs. However individual stars are not resolved, and hence ’pixel’ lensing techniques are required (see Sec. 3.5). See Refs. Kerins et al. (2001); Baltz et al. (2003) for theoretical studies of microlensing towards M31, including calculations of the differential event rate.
The standard calculation of the halo differential event rate outlined above, Eq. (10), assumes that the dark matter distribution is smooth, and fully specified by the density profile, $\rho(r)$. However PBH dark matter is expected to cluster more on sub-galactic scales than particle dark matter. Since PBHs are discrete objects, there are Poisson fluctuations in their distribution and PBH clusters start forming shortly after matter-radiation equality Afshordi et al. (2003); Inman and Ali-Haïmoud (2019). For PBHs formed from gaussian fluctuations, the clusters are sufficiently diffuse Jedamzik (2020) that the PBHs still act as individual microlenses. In this case the effect on the microlensing differential event rate, and hence constraints on the PBH mass and halo fraction, is small Petač et al. (2022); Gorton and Green (2022). Some PBHs are expected to be in binaries which could potentially lead to a source being lensed twice, however the binaries are typically sufficiently wide that the time between such repeating events is expected to be much longer that the duration of microlensing surveys Petač et al. (2022). The probability distribution of perturbations that are large enough to produce PBHs which make up a significant fraction of the dark matter, are expected to be non-gaussian, e.g. Ref. Atal and Germani (2019), however, and PBHs formed from non-gaussian fluctuations can form more compact clusters Young and Byrnes (2020). If clusters are sufficiently compact that the cluster as a whole acts as a microlens, then microlensing parameter constraints (or allowed regions) will shift to lower PBH masses Calcino et al. (2018).
2.3 Non-standard events
Deviations from the ’single point source, single point lens’ assumptions underlying the standard Paczyński light curve, Eq. (1), can have observable consequences. For a more detailed review see Ref. Mao (2012). We discuss specific observed non-standard events in Sec. 3.2.
The standard light curve is significantly modified if the impact parameter of an event is comparable to the ratio of the radius of the source star to the Einstein radius Witt and Mao (1994); Gould (1994). In particular finite source effects significantly reduce the maximum magnification for compact objects with mass $M\lesssim 10^{-9}M_{\odot}$ Sasaki et al. (2018).
The standard analysis assumes all motions are linear. However this is not the case if the lens or source are binary, and the Earth’s orbit about the Sun also violates this assumption Gould (1992). The parallax effect, due to the Earth’s orbit, is detectable if the event timescale is not significantly shorter than a year Gould (1992). For a binary source, orbital motion leads to periodic distortions in the light curve Griest and Hu (1992). This is commonly referred to as the ’xallarap’ effect (the inverse of parallax). A binary lens can generate complex, multi-peaked light curves with caustics (where the magnification formally diverges) Mao and Paczynski (1991).
PBHs with mass $M_{\rm PBH}\lesssim 10^{-11}M_{\odot}$ have Schwarzschild radius comparable to the wavelength of light, resulting in diffraction and interference effects. These wave optics effects reduce the magnification and place a lower limit on the mass of COs that can be probed by microlensing Sugiyama et al. (2020); Montero-Camacho et al. (2019).
Some of these effects can break the microlensing degeneracy mentioned above, and allow the lens mass and/or projected transverse velocity to be measured. The latter is useful as it acts as a diagnostic for the lens location Boutreux and Gould (1996). For lenses in the MW halo and LMC, the projected transverse velocity is determined by the velocity dispersion of the MW halo and LMC, which are small and large respectively. For lenses in the MW disc it is determined by the differential rotation of the Sun and disk, which has an intermediate value.
2.4 Astrometric microlensing
The above discussion focuses on photometric microlensing, i.e. the change in the apparent brightness of a source during a microlensing event. This is the observable used by the microlensing surveys to date which we will overview in Sec. 3. Astrometric microlensing is the change in the apparent position of the source Miyamoto and Yoshii (1995); Walker (1995), and in the future telescopes which accurately measure the positions of stars (astrometry) will also be able to probe CO DM using astrometric microlensing (see Sec. 4.1).
The centroid shift of the source, $\delta\theta(u)$, is given by
$\delta\theta(u)=\frac{u}{u^{2}+2}\theta_{\rm E}\,$, (17)
where $\theta_{\rm E}$ is the angular Einstein radius, $\theta_{\rm E}=R_{\rm E}/D_{\rm L}$. Using Eq. (2), $\theta_{\rm E}$ can be expressed in terms of the lens-source parallax, $\pi_{\rm rel}$,
$\pi_{\rm rel}=\pi_{\rm L}-\pi_{\rm S}=\frac{1}{D_{\rm L}}-\frac{1}{D_{\rm S}}=\frac{D_{\rm S}-D_{\rm L}}{D_{\rm L}D_{\rm S}}\,$, (18)
in milliarcseconds (mas), as
$\theta_{\rm E}^{2}=8.144\,{\rm mas}\left(\frac{M}{M_{\odot}}\right)\left(\frac{\pi_{\rm rel}}{{\rm mas}}\right)\,.$ (19)
Therefore if the parallax can also be measured, the lens mass can be determined Miyamoto and Yoshii (1995); Belokurov and Evans (2002). Astrometric microlensing signals are detectable for large lens-source separations, so the optical depth is larger than for photometric microlensing Miralda-Escude (1996); Dominik and Sahu (2000).
3 Observations
In this Section we first overview the general principles of stellar microlensing observations (Sec. 3.1), before looking at the results of specific surveys which have published constraints on the halo fraction of COs in Secs. 3.2-3.6. Unless otherwise mentioned, all limits (or parameter values) quoted assume the standard halo model, see Sec. 2.2. Sec. 3.7 contains a summary of the current observational status. Figure 4 shows the final published parameter constraints on the halo fraction, $f$, and CO mass, $M$, assuming all CO have the same mass. Table 2 summarizes the main features of the key LMC microlensing surveys, including the number of events they observed. This Section aims to provide theorists with an introduction to the general principles and current status of microlensing surveys. For further details see Ref. Moniez (2010), a 2010 review of microlensing from an observational perspective.
3.1 General principles
The first step in analyzing data from a microlensing survey is to generate light curves for individual objects. Historically this was done using point spread function photometry, e.g. Refs. Alcock et al. (2001); Ansari et al. (1996). More recently different image analysis (DIA), which is more computationally demanding, but better suited to the crowded stellar fields in the MC, has been used, e.g. Refs. Zebrun et al. (2001); Lasserre et al. (2000).
[…]
As discussed in Ref. Griest and Thomas (2005), selection criteria inevitably have to balance false positives (candidate events that are not due to microlensing) and false negatives (microlensing events that are rejected). The detection efficiency calculation takes into account the later but not the former. Event selection (and detection efficiency estimation) is complicated by the $\sim 10\%$ of microlensing events, which are ’exotic’ and don’t follow the standard Paczyński light curve produce by a single point source and lens, e.g. Ref. Bennett et al. (2005). In some cases, e.g. MACHO-LMC-9 which we discuss further in Sec. 3.2, the event is unambiguously microlensing and the features in the light curve provide information about the location of the lens. However in other cases non-microlensing variability can be fit by a non-standard lens model (see discussion in Ref. Bennett et al. (2005)). As discussed further in Sec. 3.2, in their final 5.7 year results, the MACHO collaboration used two sets of selection: criteria A designed to detect high-quality simple events, and criteria B designed to also select exotic and lower signal to noise events Alcock et al. (2000). EROS-2 Tisserand et al. (2007) and OGLE Mróz et al. (2025) carried out manual inspections of a subset of their light curves to check for a significant population of overlooked events.
The measured optical depth can be estimated using
$\tau=\frac{1}{E}\frac{\pi}{4}\sum_{i}\frac{\hat{t}_{i}}{\epsilon(\hat{t}_{i})}\,$, (20)
[…]
3.2 MACHO
Between 1992 and 1999 the MACHO collaboration monitored over 10 million stars in the LMC. Their first year results contained 3 candidate events and (assuming all 3 were due to lens in the MW halo) they found a best-fit halo fraction $f=0.2_{-0.1}^{+0.2}$ and a MACHO mass $M=0.065^{+0.06}_{-0.03}M_{\odot}$ Alcock et al. (1996). In their 2-year data they found 8 candidate events Alcock et al. (1997). MACHO-LMC-2 and 3 from their 1-year data set did not pass the revised event selection cuts used in the 2-year analysis (and one of these two events subsequently brightened again repeatedly and is therefore thought to be a variable star). On the other hand, 3 of the 8 2-year-candidates were in the 1-year data but did not pass the 1-year cuts Alcock et al. (1997). MACHO-LMC-9 had a U-shaped light curve, characteristic of a binary lens and, if the source star is single, the lens is likely to lie in the LMC Bennett et al. (1996). Using a sub-sample of 6 events (excluding the binary lens, MACHO-LMC-9, and MACHO-LMC-10 which they deemed a marginal candidate) they found a halo fraction $f=0.5_{-0.2}^{+0.3}$ and a MACHO mass $M=0.5_{-0.2}^{+0.3}M_{\odot}$. A specific search for short timescale events in the 2-year data excluded COs in the range $3\times 10^{-7}\lesssim M/M_{\odot}\lesssim 5\times 10^{-4}$ making up all of the MW halo Alcock et al. (1996). A subsequent EROS-MACHO joint analysis tightened this limit to $f<0.10$ for this planetary mass range Alcock et al. (1998).
The final MACHO results were based on 5.7-years of observations Alcock et al. (2000). They used two sets of selection criteria. Criteria A used similar cuts to their two year analysis Alcock et al. (1997) and was designed to select simple microlensing events, with a single, high significance bump in the light curve and a flat baseline. Criteria B was intentionally looser (with higher detection efficiency), and designed to search for exotic and/or low signal-to-noise events. Marginal events suspected of being supernovae in background galaxies were included in set B but rejected from set A. Set A contained 13 events (MACHO-LMC 1, 4, 5, 6, 7, 8, 13, 14, 15, 18, 21, 23 and 25) while an additional 4 events (9, 20, 22 and 27) passed criteria B  [^3] The event numbering reflects the fact that events 2 and 3 in the 1-year data were subsequently rejected, and a further 8 candidate events were classed as supernovae in background galaxies.. The best fit halo fraction (assuming a standard halo) decreased to $f\approx 0.2\pm 0.1$, with the best fit mass remaining of order $0.5M_{\odot}$. The contribution of the MW disk and LMC to the total differential event rate were taken into account, with $\sim 3$ events expected to be due to stars.
Follow-up observations have subsequently revealed further information about the nature of the lenses in some events. We summarize these developments, see also Ref. Mróz et al. (2025) for more extensive, recent discussion. Ref. Bennett et al. (2005) presented additional data for MACHO-LMC-4, 13 and 15. They found that all 3 events were well fit by a standard Paczyński light curve, supporting their microlensing interpretation. Ref. Bennett et al. (2005) also discussed the interpretation of other MACHO events. The light curve of MACHO-LMC-23 deviated from the standard Paczyński form, and EROS and OGLE both observed the star brightening again Bennett et al. (2005); Tisserand et al. (2007); Wyrzykowski et al. (2011), so it is therefore very unlikely to be a microlensing event.
MACHO-LMC-5 is a so called ’jerk-parallax’ event Drake et al. (2004), where the light curve depends on the 2nd derivative of the Earth’s velocity Gould (2004). Analysis of Hubble Space Telescope data allowed the parallax and proper motion of the lens to be measured, and the lens identified as an M dwarf star in the MW disk Drake et al. (2004), as suggested by Ref. Gould et al. (1997). High resolution follow up data from the Global Microlensing Alert Network found a periodic modulation in the light curve of MACHO-LMC-14 Alcock et al. (2001). These oscillations are thought to be the so-called ’xallarap’ effect which occurs when the source is binary Griest and Hu (1992). This allowed the lens projected velocity to be estimated, and the resulting value is more consistent with expectations for a lens located in the LMC rather than the MW Alcock et al. (2001). The light curve of MACHO-LMC-1 has an anomaly near its peak, which suggests that the lens is binary Dominik and Hirshfeld (1994, 1996); Mróz et al. (2025).
Follow-up infrared observations of MACHO-LMC-20 (which passed the B criteria) identified the lens as a M dwarf in the MW disk Kallivayalil et al. (2006). Follow-up observations of MACHO-LMC-22 (which was excluded from set A by hand) found that the source is extended and has emission lines not characteristic of stars, and is therefore now thought to be a supernova or active galactic nucleus in a background galaxy Allsman et al. (2001). After MACHO-LMC-22 is excluded, none of the MACHO LMC events have Einstein diameter crossing times longer than 150 days, which excludes CO with mass $0.3\lesssim M/M_{\odot}\lesssim 30$ making up more than $\sim 60\%$ of the MW halo Allsman et al. (2001).
Belokurov, Evans and Le Du analyzed the MACHO light curves using neutral networks and concluded that 7 of the criteria A events (MACHO-LMC-1, 5, 6, 14, 23  [^4] Event 23 has varied again and is likely a variable star, see discussion above and Refs. Bennett et al. (2005); Tisserand et al. (2007); Wyrzykowski et al. (2011) and 25) were convincing microlensing candidates Belokurov et al. (2004). Bennett also revisited the interpretation of the 13 set A MACHO events Bennett (2005). He found that 7 of the events are confirmed microlensing candidates (MACHO-LMC-1, 4, 5, 13, 14, 15 and 25) and argued that between 10 and 12 events are likely to be microlensing (with the up to 2 remaining events being variable stars), significantly larger than the $\sim 2-3$ events expected from stars in the MW or LMC, and consistent with $f\approx 0.16\pm 0.06$ of a standard MW halo being in CO. Subsequently, OGLE found that MACHO-LMC-7 (which was labeled ’unconfirmed’ by Bennett Bennett (2005)) varied repeatedly and is hence a variable star Wyrzykowski et al. (2011). OGLE also found that MACHO-LMC 8 and 18 (which were also labeled ’unconfirmed’ in Ref. Bennett (2005)) had periodic variability in their baseline Mróz et al. (2024). They argue that this is not common for microlensing events, and hence these events are likely due to variable stars, and the best fit halo fraction is in fact $\sim 40\%$ smaller than the value ($f\approx 0.2$) originally found in Ref. Alcock et al. (2000). Two of the MACHO set A events occurred during the operation of EROS-2 Tisserand et al. (2007). The star in MACHO-LMC-15 was too dim to be seen by EROS-2, while MACHO-LMC-14 was observed by EROS2, but only in one colour, and was therefore not included in their sample.
As part of their response to criticism by Hawkins and García-Bellido Hawkins and García-Bellido (2025), the OGLE collaboration have published a reanalysis of the MACHO candidate events using more modern DIA techniques Mróz et al. (2025). They deem events MACHO-LMC-6, 13 and 15 weak candidates, due to their light curves being either asymmetric, or lacking sufficient data around the peak. They note, however, that Bennett et al. Bennett et al. (2005) presented additional follow-up photometry which supported the microlensing interpretation of events 13 and 15. Based on their new DIA light curves, they conclude that 7 of the MACHO events are very likely microlensing (MACHO-LMC-1, 4, 5, 9, 14, 21, 15) and argue that the other candidate events would be unlikely to pass ’sensible microlensing cuts’.
Table  1 contains a summary of the current status of the 13 events which passed criteria A. Of the 13 events, 4 are now thought to be variable stars and 2 are microlensing events with the lens likely not in the MW halo. This leaves 6 or 7 candidate lensing events where the lens could be located in the MW halo.
The MACHO collaboration did not publish CO constraints from observations towards the SMC, but they did observe 2 candidate microlensing events: MACHO97-SMC-1 Alcock et al. (1997) (which was also detected by EROS Palanque-Delabrouille et al. (1998) and OGLE Udalski et al. (1997)) and MACHO98-SMC-1. The later was a caustic crossing binary lens event. Its live detection facilitated observations by multiple microlensing collaborations and an accurate measurement of a low value for the lens projected motion, indicating that the lens is likely in the SMC rather than the MW halo Alcock et al. (1999).
Table 1: A summary of the current status of the 13 events which passed criteria A in the 5.7 year MACHO collaboration analysis Alcock et al. (2000). See also Table 1 of Ref. Mróz et al. (2025).
[…]
Event Status
1 lens is possibly binary Dominik and Hirshfeld (1994, 1996); Mróz et al. (2025)
4
5 lens is star in MW disk Drake et al. (2004)
6
7 variable star (has varied repeatedly) Wyrzykowski et al. (2011)
8 more likely a variable star Mróz et al. (2024)
13
14 lens likely in LMC Alcock et al. (2001)
15
18 more likely a variable star Mróz et al. (2024)
21
23 brightened a 2nd time, likely a variable star Bennett et al. (2005)
25
3.3 EROS
EROS ran from 1990 to 2003. Its first phase, EROS-1 found two candidate LMC events Ansari et al. (1996); Renault et al. (1997). These events were compatible with microlensing by COs with mass $\approx 0.1M_{\odot}$. However, because of the unusual nature of these events (the source star for one was an eclipsing binary star Ansari et al. (1996)), and the possibility of having detected variable stars, they chose to place an upper limits on the halo fraction: $f<0.2$ for $10^{-7}\lesssim M/M_{\odot}\lesssim 0.01$. A dedicated search for short-timescale events excluded planetary mass COs, with mass $10^{-7}\lesssim M/M_{\odot}\lesssim 10^{-4}$, making up more than $10\%$ of the MW halo Renault et al. (1998).
In their first two years of SMC observations, EROS-2 found one candidate event, EROS2-SMC-1 Palanque-Delabrouille et al. (1998); Afonso et al. (1999), also detected by MACHO (MACHO97-SMC-1) Alcock et al. (1997) and OGLE Udalski et al. (1997). In the first 2 years of EROS-2 LMC observations they found 2 events (EROS2-LMC-3 and 4). Combined with their surviving previous candidate events (EROS-LMC-2 was eliminated having undergone a 2nd variation Lasserre et al. (2000)) they excluded COs with $10^{-7}(10^{-6})\lesssim M/M_{\odot}\lesssim 1(10^{-3})$ making up more than $40(10)\%$ of the MW halo Lasserre et al. (2000).
With another three years of SMC observations EROS-2 found an additional three events (EROS2-SMC-2-4), with Einstein radius crossing times >200>200 days (longer than the events observed towards the LMC) Afonso et al. (2003). They argued that these events "seem more compatible with unidentified variable stars or self-lensing within the cloud than with halo objects". This was partly based on a prior expectation that halo COs would not have super-Solar masses, however follow-up observations concluded that these events were indeed long-period variables Tisserand et al. (2007). The SMC observations placed slightly weaker constraints on the halo fraction than the previous LMC observations.
The final EROS-2 analysis used observations of a subsample of 7 million bright stars in the MC, in order to aid discrimination of variable stars, and minimize the effects of blending Tisserand et al. (2007). They found only one event, and constrained COs with $10^{-6}\lesssim M/M_{\odot}\lesssim 1$ to make up less than $10\%$ of the MW halo. Section 6 and Table 4 of Ref. Tisserand et al. (2007) contains an update on the status of the 11 previously published EROS MC events. Only 1 (EROS-SMC-1) remains a microlensing candidate. EROS1-LMC-1 and 2, and EROS2-LMC-3 and 4 underwent a 2nd variation. With improved data analysis EROS2-LMC-5, 6 and 7 were classified as Supernovae, and (as mentioned above) additional data showed that EROS2-SMC-2, 3 and 4 were long period variable stars.
Ref. Blaineau et al. (2022) combined archival data from EROS-2 and MACHO to probe long timescale events, and concluded that CO with mass $10\lesssim M/M_{\odot}\lesssim 100(10^{3})$ have $f<0.15(0.5)$. Combined with the previous EROS results, $f>0.15$ was excluded for $10^{-6}\lesssim M/M_{\odot}\lesssim 100$.
3.4 OGLE
The OGLE project has run for more than thirty year, starting in 1992 Soszyński (2024). The first OGLE MC observations (OGLE-II) found two candidate LMC events Wyrzykowski et al. (2009). They found that if both were due to COs the halo fraction was $f=0.08\pm 0.06$. However they argued that these events were consistent with expectations from self-lensing, and constrained CO with $0.01\lesssim M/M_{\odot}\lesssim 0.2$ to make up less than 10% of the MW halo. OGLE-LMC-02 subsequently underwent multiple additional brightenings and is therefore not a microlensing event Mróz et al. (2024). Their initial observations towards the SMC found a single, unconvincing, candidate and slightly weaker limits Wyrzykowski et al. (2010).
In 8 years of observations towards the SMC, OGLE-III found 3 convincing events (OGLE-SMC-2-4) Wyrzykowski et al. (2011). The light curve of OGLE-SMC-2 exhibits a clear parallax effect, and also deviates from the simple Paczyński curve around its peak. Follow up observations (including data from the Spitzer space telescope) measured the lens projected velocity and found that the lens is most likely a BH binary in the MW halo, with mass $M\sim 10M_{\odot}$ Dong et al. (2007). Ref. Wyrzykowski et al. (2011) argued that this event is consistent with BHs making up of order a per-cent of the MW halo. Like MACHO-LMC-5, OGLE-SMC-3 was a jerk-parallax event, with the lens likely a M dwarf in the MW disc Wyrzykowski et al. (2011). The combined OGLE-II and III MC limits on the halo fraction were similar to those from EROS2 for $M\gtrsim 0.01M_{\odot}$ Wyrzykowski et al. (2011).
Results from OGLE-III and IV observations of 78.7 million stars in the LMC for up to 20 years were published in 2024 Mróz et al. (2024). Their automated data pipeline found 13 events, all with Einstein radius crossing times less than one year. A manual, ’by eye’, search found an additional 3 events. They found that the number of events, and their positions, timescales and parallaxes, are consistent with expectations from stars in the MW disk and the LMC. Using the 13 events identified using the automated data pipeline, they found that COs in the mass range $10^{-4}(10^{-5})\lesssim M/M_{\odot}\lesssim 6(900)$ can not make up more than $1(10)\%$ of the MW halo (assuming the MW halo has a contracted Navarro Frenk White density profile fitted to Gaia DR2 and other data in Ref. Cautun et al. (2020)) Mróz et al. (2024). There are also similar, but slightly weaker, constraints from OGLE’s 20 years of SMC observations, with the six observed events being consistent with expectations from stars in the SMC or MW disc Mróz et al. (2025).
Niikura et al. Niikura et al. (2019) used OGLE microlensing observations towards the Galactic bulge to probe planetary mass COs. They found that most events were well-fit by microlensing by stellar populations, and excluded $f>0.1$ for $10^{-6}\lesssim M/M_{\odot}\lesssim 10^{-2}$. There was a distinct population of six ultra-short ($\sim 0.1$ day) events. These events could be due to free-floating planets, however they argued that these events are also consistent with Earth mass PBHs making up a few per-cent of the MW halo. From 2022-2024 OGLE carried out a high-cadence survey of the MC, observing 35 million stars up to 21 times per night Mróz et al. (2024). They only found one event, with Einstein radius crossing time $\sim 90$ days, which allowed them to exclude COs with mass $10^{-8}\lesssim M/M_{\odot}\lesssim 0.01$ making up more than $1\%$ of the MW halo.
3.5 M31
Various collaborations have carried out so-called pixel microlensing surveys towards M31. Stars in M31 are too distant to be resolved individually, so instead microlensing surveys search for variations in the flux of the pixels of the image. Ref. Calchi Novati (2010) contains a review of pixel microlensing surveys towards M31 as of 2009.
In the early 2000s POINT-AGAPE Calchi Novati et al. (2005) found 6 events, significantly more than the $\sim 1$ they expected from self-lensing. They concluded that if the average mass of COs is $\sim 1(0.01)M_{\odot}$, then at 95% confidence they make up more than $20(8)\%$ of the MW halo. Subsequently, MEGA de Jong et al. (2006) and PLAN Calchi Novati et al. (2014) both concluded (using detailed calculations of the detection efficiency, and also more sophisticated mass models of M31 to calculate the self-lensing rate) that their observed event rates were broadly consistent with self lensing, and constrained the MW halo fraction in Solar mass COs to be less than $(20-30)\%$.
More recently high cadence (2 minute sampling interval) observations of M31 have been carried out using the Hyper Suprime-Cam (HSC) on the Subaru telescope Niikura et al. (2019); Sugiyama et al. (2026). Initial observations found one candidate event which led to a constraint $f\lesssim 10^{-2}$ for $10^{-5}\lesssim M/M_{\odot}\lesssim 10^{-9}$ Niikura et al. (2019, 2019). The finite size of the lensed stars reduces the magnification Witt and Mao (1994), and makes the constraint weaker than it would otherwise be, for $M\lesssim 10^{-7}M_{\odot}$ Niikura et al. (2019); Smyth et al. (2020). Very recently this work has been updated with additional data, and improved analysis of the original data (in particular the inclusion of finite source size effects in the detection efficiency estimation) Sugiyama et al. (2026). Twelve microlensing candidates have been identified, with 4 being classed as secure, high-significance, detections. If these events are due to COs in the MW halo the corresponding mass and halo fraction are $M\sim(10^{-7}-10^{-6})M_{\odot}$ and $f\sim 0.01-0.1$.
3.6 Kepler
In the early 2010s Griest and collaborations searched for short timescale microlensing events in publicly available observations of nearby stars (designed to search for extra-Solar planets) by Kepler. Their final results Griest et al. (2014), based on 2 years of data, excluded compact objects with $10^{-9}\lesssim M/M_{\odot}\lesssim 10^{-7}$ making up all of the MW halo.
3.7 Summary
[…]
Figure 4: Constraints on the CO mass, $M$, and MW halo fraction, $f$, at 95% confidence, assuming all CO have the same mass. The solid lines show the exclusion limits from EROS (’E’) Tisserand et al. (2007) (yellow), MACHO long timescale (’M-long’) Allsman et al. (2001) (red), EROS plus MACHO long timescale (’E+M’) Blaineau et al. (2022) (orange), OGLE Mróz et al. (2024) and OGLE high cadence (’OGLE-hc’) Mróz et al. (2024) (blue), and Kepler (’K’) Griest et al. (2014) (green). The purple lines are from high cadence observations of M31 using the Subaru-HSC. ’HSC19’ shows the original exclusion limit Niikura et al. (2019) as calculated in Ref. Croon et al. (2020), while ’HSC26’ shows the exclusion limit from the latest analysis, under the assumption that none of the events observed are due to CO in the MW halo. The dashed line shows the allowed parameter values assuming that the 4 ’secure’ high-significance events are due to CO in the MW halo. The dashed red line shows the allowed parameter values from the 13 set A events in the final 5.7 year MACHO data set, as published in Ref. Alcock et al. (2000). Subsequent observations indicate that some of these events are not microlensing (see Sec. 3.2 and Refs. Tisserand et al. (2007); Mróz et al. (2025) for discussion). The dotted blue line, labeled ’OGLE?’, shows the parameter values found assuming that the ultra-short duration events observed towards the Galactic bulge by OGLE are due to COs in the MW halo Niikura et al. (2019). The dot-dashed black line shows the envelope of the tightest constraint for each value of $M$. This plot was made using a modified version of Kavanagh’s PBHBounds Kavanagh .
Figure 2 shows the final published parameter constraints on the halo fraction, $f$, and compact object mass, $M$, (assuming all compact objects have the same mass) from the surveys discussed in this section  [^5] Results from M31 surveys, other than HSC, are not shown. These surveys quoted approximate allowed or excluded values for the halo fraction, but they did not publish exclusion limit curves or allowed regions of parameter space.. Table 2 contains a summary of the key features of the MACHO Alcock et al. (2000), EROS Tisserand et al. (2007) and OGLE Mróz et al. (2024) LMC microlensing surveys. See Table 1 of Ref. Moniez (2010) for a more detailed comparison of stellar microlensing surveys (as of 2010).
Table 2: A summary of the main features of the MACHO Alcock et al. (2000), EROS Tisserand et al. (2007) and OGLE Mróz et al. (2024) LMC microlensing surveys. The 13 MACHO events are from their tighter selection criteria A. Four of these events have subsequently been found to likely be variable stars, see Ref. Tisserand et al. (2007); Mróz et al. (2025) and Sec. 3.2. The 13 events listed for OGLE are those found using their automated search pipeline, and used for their halo fraction constraints, see Sec. 3.4.
[…]
Survey Duration Field of view Number of stars, $N_{\star}$ Number of events, $N_{\rm obs}$
(years) (${\rm deg}^{2}$) (million)
MACHO 5.7 14 12 13
EROS $\sim$ 12 88 (EROS-2) $\sim$ 25 (EROS-2) 1
OGLE $\sim$ 20 105 (OGLE-IV) 78.7 (OGLE-IV) 13
4 Future
4.1 Observations
In the near future new telescopes will be able to carry out next generation microlensing surveys. The Nancy Grace Roman Space Telescope (formerly known as WFIRST) has excellent resolution, which will reduce the effects of blending and allow the measurement of finite source effects. It will carry out high cadence observations, as required to detect short timescale events and probe light CO. The ground-based Vera C. Rubin Observatory (formerly known as the Large Synoptic Survey Telescope) will carry out a long duration wide-field survey. It will have high sensitivity to long timescale events, and hence intermediate and high mass CO. Telescopes with high spatial resolution, capable of high-precision astrometry, will be able to carry out astrometric microlensing measurements and probe the angular Einstein radius (see Sec. 2.4).
Following early work in Ref. Cieplak and Griest (2013), Ref. DeRocco et al. (2024) showed that the Roman Space Telescope’s Galactic Bulge Time Domain Survey will be able to to discriminate planetary-mass PBHs from free floating planets, and probe halo fractions $f\sim 10^{-3}-10^{-4}$ using the timescales of (photometric) microlensing events. Ref. Fardeen et al. (2024) showed that this survey could probe $f\sim 10^{-3}$ for $10^{-1}\lesssim M/M_{\odot}\lesssim 10^{2}$ if it is possible to distinguish PBHs and stellar lenses using the joint distributions of the timescales and angular Einstein radii of events. If it is not possible to do this then the sensitivity will be reduced to $f\sim 10^{-1}$, weaker than existing OGLE limits in this mass range Mróz et al. (2024). Similar constraints may also be possible using the Rubin Observatory Drlica-Wagner et al. (2019). With improved understanding of backgrounds, Gaia time-series data could probe halo fractions down to $f\sim 10^{-3}$ for multi-Solar mass PBHs Verma and Rentala (2023), see also Ref. Belokurov and Evans (2002). Thirty-metre class, Extremely Large Telescopes, will also be able to probe PBHs via astrometric microlensing Bird et al. (2023). Combining photometric and astrometric microlensing data would aid the detection and characterization of events Pruett et al. (2024). In all cases, correctly identifying events and minimizing the false positive rate will be crucial (as we have seen for previous surveys in Sec. 3). See Ref. Crispim Romao et al. (2025) for work in this direction.
4.2 Theory
The key theoretical uncertainty is the distribution of PBHs, specifically their local velocity distribution and also whether PBHs form compact clusters. As discussed in Sec. 2.2, uncertainties in the DM distribution lead to uncertainties in the differential event rate and hence the constraints on the halo fraction. If the DM is smoothly distributed, the most significant uncertainty is the dependence of the rate of short-timescale events, and hence the constraints on $f$ for large masses, on the local velocity distribution Green (2025). If PBHs form compact clusters, such that the cluster as a whole lenses, the constraints would change significantly Calcino et al. (2018). Understanding the present day spatial distribution of PBHs formed from realistic density perturbation distributions (i.e. how many PBHs are in clusters, and what are their properties? how many PBHs are in binaries, and what are their properties?) is a major outstanding challenge.
5 Summary
We have explored how stellar microlensing surveys probe compact object dark matter, in particular primordial black holes. Early indications from the MACHO collaboration that a significant fraction of the Milky Way halo could be in the form of roughly Solar Mass COs Alcock et al. (1997, 2000) were not confirmed by EROS Tisserand et al. (2007) and OGLE Mróz et al. (2024). Long duration Mróz et al. (2024) and high cadence Niikura et al. (2019); Mróz et al. (2024) observations have extended the range of CO masses probed by microlensing. Under the standard assumptions outlined in Sec. 2, collectively these surveys exclude COs with mass $10^{-11}\lesssim M/M_{\odot}\lesssim 10^{4}$ making up all of the DM in the MW, while those in the mass range $10^{-10}\lesssim M/M_{\odot}\lesssim 1$ can make up less than one percent. Further theoretical work is required to ascertain how robust the standard assumptions are, in particular regarding the spatial distribution of PBHs. Future observations, in particular astrometric ones using the Roman Telescope and Rubin Observatory, can potentially probe PBHs down to $f\sim 10^{-3}-10^{-4}$. However further work is required to develop techniques for discriminating PBH-induced lensing events from backgrounds (including, but not limited to, lensing events due to astrophysical objects).
[…]
This research was supported by an STFC Consolidated Grant [Grant No. ST/T000732/1] and a Leverhulme Trust Research Fellowship [Grant No. RF-2025-282].
[…]
No new data were created or analyzed in this study. Data sharing is not applicable to this article.
[…]
The author declares no conflicts of interest.
[…]
Abbreviations The following abbreviations are used in this manuscript:
CO Compact Object
DM Dark Matter
DIA Difference Image Analysis
LMC Large Magellanic Cloud
MC Magellanic Cloud
MW Milky Way
PBH Primordial Black Hole
SMC Small Magellanic Cloud
[…]
no
[…]